OpenStudy (anonymous):

Find the slope of the line through (3, –6) and (5, 8).

OpenStudy (anonymous):

can u help me

OpenStudy (zepp):

The slope(m) could be found using m = dy/dx

OpenStudy (zepp):

dy = y2-y1 dx = x2-x1 (3(x1), –6(y1)) and (5(x2), 8(y2)). Good luck ;D

OpenStudy (anonymous):

7?

OpenStudy (zepp):

y2-y1 = 8- -6 = 14 x2-x2 = 5 - 3 = 7 m = dy/dx m = 14/7 m = 2 :D